OpenStudy (wildflower):

Does the data in this table represent a function? Input (x) 3 7 11 15 Output (y) 4 6 8 10

satellite73 (satellite73):

as long as the first numbers are all different, the answer is YES a function ignore the y's completely

OpenStudy (wildflower):

what would it look like as a function?

satellite73 (satellite73):

just as you see it, the table

satellite73 (satellite73):

you could also make ordered pairs \[\{3,4),(7,6),(11,8),(15,10)\}\]

OpenStudy (wildflower):

f(x) = 3x + 4 ? is this right?

satellite73 (satellite73):

not hardly

satellite73 (satellite73):

if it was \(f(x)=3x+4\) then the y value would be 4 more than 3 times the x valuie

satellite73 (satellite73):

so here you have \((3,4)\) as pair, that means \[f(3)=4\]

OpenStudy (wildflower):

i know there this a difference of 4 in the x outputs

satellite73 (satellite73):

looks to me like each output increases by 2, not by 4

satellite73 (satellite73):

it is the inputs that are increasing by 4 each time

satellite73 (satellite73):

lets back up a second this is the entire question Does the data in this table represent a function? right?

OpenStudy (wildflower):

oh yeah sorry, its the x inputs

OpenStudy (wildflower):

there are 3 parts to the question, this is the first part

OpenStudy (wildflower):

in the next part they ask to "Compare the data in the table with the relation f(x) = 5x – 21. Which relation has a greater value when x = 11?"

satellite73 (satellite73):

ok well that part only asks "function, yes/no" it does not ask you to find a rule for the function all you need to see is that all the x values are different, then you answer YES

OpenStudy (wildflower):

ok

OpenStudy (wildflower):

in part A, they asked me to justify my answer.

satellite73 (satellite73):

ok if \[f(x)=5x-21\] do you know how to find \(f(11)\)?

OpenStudy (wildflower):

no

satellite73 (satellite73):

justify your answer "all the x values are distinct (different)"

satellite73 (satellite73):

ok "no" is a fine answer

satellite73 (satellite73):

\[f(x)=5x-21\\ f(\xi)=5\xi -21\\ f(\spadesuit)=5\spadesuit-21\\ f(11)=?\]

OpenStudy (wildflower):

I don't know.

satellite73 (satellite73):

\[f(11)=5\times 11-21=?\]

OpenStudy (wildflower):

55-21 =34

satellite73 (satellite73):

yes that is what i get too

satellite73 (satellite73):

and what is the y value from the table that corresponds to the x value of 11?

OpenStudy (wildflower):

ok thank you.

OpenStudy (wildflower):

8?

satellite73 (satellite73):

yes

satellite73 (satellite73):

now you can answer in the next part they ask to "Compare the data in the table with the relation f(x) = 5x – 21. Which relation has a greater value when x = 11?" easily enough, since one output is 8 (from the table) and the other output is 34

OpenStudy (wildflower):

they ask to turn the data in the table into a function, and compare it to f(x) = 5x - 21. I don't know what the function is from the data in the table. I need help to figure that out

satellite73 (satellite73):

aren't the ordered pairs good enough, or do you need a rule?

OpenStudy (wildflower):

i think i need a function like f(x) = 5x - 21 to compare it to

satellite73 (satellite73):

ordered pairs look like\[\{(3,4),(7,6),(11,8),(15,10)\}\] rule looks like \[g(x)=sommat\]

OpenStudy (wildflower):

but i don't know how to make a function from that table

satellite73 (satellite73):

ok that is going to require some work

satellite73 (satellite73):

you already said the x values increase each time by 4 right?

satellite73 (satellite73):

and i said the y values increase by 2 that means this is a line, and the slope of the line that goes "right 4, up 2" is \[\large \frac{2}{4}=\frac{1}{2}\]

OpenStudy (wildflower):

yes

satellite73 (satellite73):

so it is going to look like \[g(x)=\frac{1}{2}x+b\] now you need \(b\)

OpenStudy (wildflower):

ok so that is the slope?

satellite73 (satellite73):

yes

satellite73 (satellite73):

right 4 , up 2 slope is 2/4

OpenStudy (wildflower):

ok

OpenStudy (wildflower):

but i think it has to start like this: f(x) = 1/2 + or - another number.

satellite73 (satellite73):

it is going to be \[g(x)=\frac{1}{2}x+ number\]

OpenStudy (wildflower):

okay this is part b: Part B: Compare the data in the table with the relation f(x) = 5x – 21. Which relation has a greater value when x = 11? so far, i did this: f(x) = 5x - 21 f(11) = 5 x 11 - 21 f(11) = 55 - 21 = 34. f(11) = 34.

OpenStudy (wildflower):

i think now i need to make a relation from the table and compare it to f(11)= 34.

satellite73 (satellite73):

no you do not

OpenStudy (wildflower):

i don't know what to do

satellite73 (satellite73):

in the table, the y value for 11 is 8 in the function the y value for 11 is 34 they ask you which is bigger i pick 34, you?

OpenStudy (wildflower):

oh okay. i think i understood. so how to i show my work?

satellite73 (satellite73):

ikd just say it you already showed it for \(f(11)=34\) there is nothing to compute for the table, it is just given to you

OpenStudy (wildflower):

is this okay for the first part? f(x) = 5x - 21 f(11) = 5 x 11 - 21 f(11) = 55 - 21 = 34. f(11) = 34.

satellite73 (satellite73):

yes for \(f(11\) sure

OpenStudy (wildflower):

is this okay too? The y-value for 11 is 8. When you compare 34 to 8, 34 is greater in value.

satellite73 (satellite73):

yes

OpenStudy (wildflower):

okay, sorry for bothering you. I have one more part left! Part C: Using the relation in Part B, what is the value of x if f(x) = 99?

satellite73 (satellite73):

solve \[5x-21=99\]for \(x\)

OpenStudy (wildflower):

okay

satellite73 (satellite73):

should take two steps only

OpenStudy (wildflower):

am i doing it wrong? 5x - 21 = 99 5x - 21 + 5 = 99 + 5 26 + 104 = 130 130 ÷ 5 = 26

OpenStudy (wildflower):

x = 26 ??

satellite73 (satellite73):

nope

satellite73 (satellite73):

add 21, divide by 5

OpenStudy (wildflower):

where?

satellite73 (satellite73):

\[5x-21=99\] add \(21\) to both sides

OpenStudy (wildflower):

so 5x - 21 + 21 = 99 +21 42 + 120 = 162 162 ÷ 5 = 32.4 I don't think thats right.

satellite73 (satellite73):

no it is not

OpenStudy (wildflower):

the answer should be 24. but i don't know how to get that

satellite73 (satellite73):

x is not a number, it is variable you do not find it until the end

satellite73 (satellite73):

\[5x - 21 + 21 = 99 +21 \] is right

satellite73 (satellite73):

but the next step should be \[5x=120\]

satellite73 (satellite73):

then divide by 5

OpenStudy (wildflower):

ok.thanks

OpenStudy (wildflower):

thank you so much for your help!!. I very much appreciate it.

OpenStudy (wildflower):

you helped a lot!